OpenStudy (anonymous):

Solve: 3x - x + 2 = 12 Answer 4 7 5 -5

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

2x+2=12 =>2x=10 =>x=5

OpenStudy (anonymous):

15-3=12

OpenStudy (ars22):

2x+2=12 2x=12-2 2x=10 x=5